OpenStudy (anonymous):

. Select the set of equations that represents the following situation: The sum of two numbers is 73, and their difference is thirty-two less than the smaller number. What are the two numbers

OpenStudy (anonymous):

x+y = 73 32 less than the smaller number???????

OpenStudy (anonymous):

x + y = 73 x-y = y-32 solving this: x= -32 +2y y +2y -32 = 73 3y = 73+32 (=) y = 35 x = 38

OpenStudy (anonymous):

y = smaller number.. well i have the same question as jimmyrep.. i didnt get the "32 less than the smaller number"

OpenStudy (anonymous):

right - im really tired at the moment and that's confused me to heck!!!1

OpenStudy (anonymous):

In my point of view what I did is the answer but i have serious doubts about the "32 less than the smaller number" thing

OpenStudy (anonymous):

97.8......BAMB THE future of rock and roll. 97.8......BAMB THE future of rock and roll.

OpenStudy (anonymous):

music!!!!!!

OpenStudy (anonymous):

diogo - you are right - their difference is 38-35 = 3 and this is 32 less than smaller number(35). i just could'nt get my head around that! i'm off to bed!!

OpenStudy (anonymous):

oh ok. Good night then